Circular No. 6974

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998dd IN ANONYMOUS GALAXY J. Maza, Department of Astronomy, University of Chile, reports the

discovery by R. Antezana, on a T-Max 400 film taken by L. E. González with the Maksutov telescope at Cerro El Roble on July 18.058 UT, of an apparent supernova (B ∼ 17) located at α = 13^{h}47^{m}33^{s}.95, δ = −33^{o}02′31′′.4 (equinox 2000.0), which is 2′′.8 east and 2′′.5 north of the nucleus of a ringed spiral galaxy. SN 1998dd was confirmed on a film taken by González using the same telescope on July 19.133. No star is visible at this position on a film obtained on June 18.139 (B > 20).

COMET C/1998 M6 (LINEAR) Improved parabolic orbital elements from MPEC 1998-O05:

T = 1998 Oct. 15.888 TT  ω = 10.^{o}047
Ω = 306.603 }2000.0
q = 5.96988 AU  i = 91.554
1998 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
July 16 20^{h}44.^{m}95 −16^{o}27.^{′}0 5.021 6.005 164.^{o}0 2.^{o}7 18.3
26 20 35.94 −15 50.2 4.985 5.998 174.8 0.9 18.3
Aug. 5 20 26.83 −15 13.1 4.985 5.991 171.9 1.4 18.3
15 20 18.02 −14 35.6 5.020 5.986 160.8 3.2 18.3
25 20 09.88 −13 57.8 5.089 5.981 149.4 4.9 18.3
Sept. 4 20 02.71 −13 20.0 5.187 5.977 138.2 6.5 18.3

COMET C/1998 M2 (LINEAR) Ephemeris extension to IAUC 6949 (elements MPEC 1998-M31):

1998 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
July 16 17^{h}42.^{m}81 +12^{o}49.^{′}8 1.926 2.742 135.^{o}2 15.^{o}1 13.8
26 17 34.30 +15 04.5 2.005 2.733 126.4 17.4 13.9
Aug. 5 17 28.29 +16 44.6 2.101 2.728 118.0 19.2 14.0
15 17 25.02 +17 56.2 2.209 2.727 110.0 20.4 14.1
25 17 24.48 +18 45.8 2.324 2.730 102.6 21.2 14.2
Sept. 4 17 26.56 +19 20.2 2.442 2.736  95.9 21.5 14.3

Visual m_{1} estimates: June 28.29 UT, 13.4 (A. Hale, Cloudcroft, NM, 0.41-m refl.); July 19.90, 14.0 (K. Hornoch, Lelekovice, Czech Rep., 0.35-m refl.).

1998 July 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green